OpenStudy (anonymous):

x and y vary inversely, and y = 2 when x = 5. Find y when x = 3. A. 0 B. 6 5 C. 10 3 D. 6

OpenStudy (aripotta):

so the choices are 0, 65, 103, or 6?

OpenStudy (anonymous):

si muchacho

OpenStudy (anonymous):

......yes

OpenStudy (aripotta):

y = k/x

OpenStudy (aripotta):

2 = k/5 can you solve for k?

OpenStudy (anonymous):

yea 10

OpenStudy (aripotta):

sorry lol. y = 10/x

OpenStudy (anonymous):

ooooh okaay

OpenStudy (aripotta):

y = 10/3

OpenStudy (anonymous):

okayyyy

OpenStudy (aripotta):

i'm guessing 103 is actually 10/3?

OpenStudy (anonymous):

okay so c?

OpenStudy (aripotta):

i guess so

OpenStudy (anonymous):

aiqht good enuf lol

OpenStudy (anonymous):

the answer choice (c) 10 3 is most likely supposed to read "10/3" but the fraction line didn't paste in well.

OpenStudy (anonymous):

so, yeah, (c)